OpenStudy (anitaasks):

How does x^2/(x^2+1) become 1-1/(x^2+1)?

OpenStudy (anonymous):

\[\frac{x^2}{x^2+1}=\frac{x^2+1-1}{x^2+1}=\frac{x^2+1}{x^2+1}-\frac{1}{x^2+1}\]

OpenStudy (misty1212):

a much easier trick than actually doing the long division fer sure

OpenStudy (misty1212):

although you would get it in one step in this case

OpenStudy (anitaasks):

Thanks!